OpenStudy (anonymous):

Evaluate the definite integral. I am using the equation writer to enter the integral below.

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\frac{ x }{ \sqrt{1+2x} }dx\]

OpenStudy (hba):

you know how to integrate ?

OpenStudy (anonymous):

yes this particular question is under the heading "The substitution rule" and is my last question for this section but I do not see a place to use the substitution rule and I do not think it has symmetry like some of the questions before it.

OpenStudy (hba):

Ok its gotta be easy then let y=1+2x and diffrentiate it

OpenStudy (anonymous):

then how do I deal with the x in the numerator? because the dy/dx piece for what you posted would just be 2 wouldn't it?

OpenStudy (hba):

I am trying to remember,what i can...

OpenStudy (anonymous):

It's often a good idea to substitute terms that are in the denominator, or under a root signs just as @hba has done. Doing this may make more terms in the numerator but those can be split up into separate fractions.

OpenStudy (hba):

@Mathmuse It's been a long time since I've done this but i am gonna try to help,I just want you to be here,so you can watch if i go wrong.

OpenStudy (anonymous):

Okay but how this is the step that I do not understand. I cannot see how you get rid of the numerator through substitution.

OpenStudy (anonymous):

I appreciate the help though

OpenStudy (hba):

Just diffrentiate it and tell me what you get ?

OpenStudy (anonymous):

The y=1+2x equation? I get that y'=2

OpenStudy (hba):

dy=2dx

OpenStudy (hba):

so, \[4\int\limits_{0}^{4}\frac{ dy }{ y }\]

OpenStudy (hba):

Looks cool to me :D

OpenStudy (anonymous):

Okay I did not do that because that is not how my textbook taught me substitution

OpenStudy (hba):

Now integrate.

OpenStudy (hba):

y'=dy/dx=2 so dy=2dx

OpenStudy (hba):

^ thats how i got it.

OpenStudy (anonymous):

i better show some steps

OpenStudy (anonymous):

isn't it \[\int\limits_{0}^{4}\frac{ 2x \times dy }{ y}\]

OpenStudy (anonymous):

I guess I am just missing what happened to the x somehow

OpenStudy (hba):

@Mathmuse Please show some step i also wanna revise.

OpenStudy (anonymous):

I am sorry but I just do not see it

OpenStudy (anonymous):

@hba I hate to bring it up but in that discussion yesterday I was saying why do her teachers need to complicate things not you guys. She did not understand me, unfortunately I was not around to correct the misunderstanding.

OpenStudy (hba):

@whatisthequestion I do not even remember what your'e talking about,link pleasee

OpenStudy (anonymous):

Starting with \[\int\limits\limits_{0}^{4}\frac{ x }{ \sqrt{1+2x} }dx\] If you want to substitute, you must get rid of any terms that are related to 'x' so we start with: \[u=1+2x\] which leads to: \[\int\limits\limits_{}^{}\frac{ x }{ \sqrt{u} }dx\] At this point there is still a dx left, so lets take the derivative of our substitution to solve for dx so we can replace it. \[u = 1+2x\] \[\frac{du}{dx} = 2\] \[\frac{du}{2} = dx\] Now we have: \[\int\limits\limits_{}^{}\frac{ x }{ \sqrt{u} }\frac{1}{2}du\]

OpenStudy (hba):

I think i deal with many problems like this daily but we have to refrain users from cheating.

OpenStudy (anonymous):

the x term in the numerator needs to go so lets rearrange our original substitution to find what x is equal to: \[u=1+2x\] \[2x=u-1\] \[x=\frac{u-1}{2}\] making this substitution and pulling the constants in front of the integral we get: \[\frac{1}{2}\frac{1}{2}\int\limits\limits_{}^{}\frac{ u-1 }{ \sqrt{u} }du\]

OpenStudy (anonymous):

Okay I never even though of going back to rewrite the 2x+1=u equation

OpenStudy (anonymous):

I just wanted to make it clear that I was not saying that you were complicating things but that her teachers were

OpenStudy (anonymous):

now we can clean it up: \[\frac{1}{4}[\int\limits\limits\limits_{}^{}\frac{ u }{ \sqrt{u} }du - \int\limits\limits\limits_{}^{}\frac{ 1 }{ \sqrt{u} }du]\] \[\frac{1}{4}[\int\limits\limits\limits_{}^{}u^{\frac{1}{2}}du - \int\limits\limits\limits_{}^{}u^{-\frac{1}{2}}du]\]

OpenStudy (hba):

@whatisthequestion Yeah don't worry i know you were helping :) @Mathmuse Let him do some work lol :D

OpenStudy (anonymous):

and I have the formulas for how to differentiate those thank you besides my final (and studying for it) that was my last math question for calculus this semester and I just could not see how to solve it

OpenStudy (anonymous):

don't forget the bounds of integration. They used to go from 0 to 4 in the x domain, but now you're taking the integral in the u domain. you have to sub those out too.

OpenStudy (hba):

Yeah you have to sub them back after your'e done .

OpenStudy (anonymous):

yes with G(x) which is =y

OpenStudy (anonymous):

expect at those points

OpenStudy (anonymous):

you don't have to sub anything after you're done. the definite integral gives you a strictly numeric answer. once you've calculated it your finished. For this question I would follow only what I've written. there is no y, there is only u.

OpenStudy (anonymous):

sorry my textbook always uses the notation u for the piece they sub in, I accidentally slipped back into that notation

OpenStudy (anonymous):

y is normally reserved for the dependent variable of a function. you will see u's and v's for substitutions and other calculus methods